# Investigating the interaction of linguistic and mathematical reasoning in language models using multilingual number puzzles

### v1

@@ Abstract
Across languages, numeral systems vary widely in how they construct and combine numbers. While humans consistently learn to navigate this diversity, large language models (LLMs) struggle with linguistic-mathematical puzzles involving cross-linguistic numeral systems, which humans can learn to solve successfully. We investigate why this task is difficult for LLMs through a series of experiments that untangle the linguistic and mathematical aspects of numbers in language. Our experiments establish that models cannot consistently solve such problems unless the mathematical operations in the problems are explicitly marked using known symbols ($+$, $\times$, etc, as in “twenty + three”). In further ablation studies, we probe how individual parameters of numeral construction and combination affect performance. While humans use their linguistic understanding of numbers to make inferences about the implicit compositional structure of numerals, LLMs seem to lack this notion of implicit numeral structure. We conclude that the ability to flexibly infer compositional rules from implicit patterns in human-scale data remains an open challenge for current reasoning models.
@@
Antara Raaghavi Bhattacharya1, 2, Isabel Papadimitriou3, Kathryn Davidson2, David Alvarez-Melis1,3 1Computer Science Department, Harvard University,  2Linguistics Department, Harvard University, 3Kempner Institute for the Study of Natural and Artificial Intelligence at Harvard University Correspondence: antarabhattacharya@college.harvard.edu
@@ 1 Introduction
Figure 1: Making operators explicit significantly improves performance. Results for explicit operator experiments, for the single-character variable case (For the results on multi-character variables, see Appendix B Figure 7). Making operators explicit shows performance improvement over the implicit condition, but this is only substantially and reliably the case when the operator is made explicit with a familiar symbol like “+”. Error bars = standard error of the mean. 10 problems, 5 iterations per problem.
@@ 2 Methods
Models. We used OpenAI o1-mini Jaech et al. (2024) and DeepSeek-R1-distill-Qwen-7B Guo et al. (2025) reasoning models to conduct our experiments, querying o1-mini via the API and running DeepSeek locally. We will publicly release all code and data used for our experiments.
Data. We obtained data for linguistics olympiad problems from two publicly available datasets: LingOly Bean et al. (2024) and Linguini Sánchez et al. (2024), filtering both datasets for problems tagged as “number systems”. After filtering, we had 15 problems from the LingOly and 8 problems from the Linguini dataset. Not every problem in the dataset could be standardized in the ways that our experiments required. The entire dataset was thus manually evaluated for suitable problems, and 10 problems were chosen for evaluation, all in distinct languages (see Appendix D). These problems spanned a range of difficulty from the first round of the UK Linguistics Olympiad to the International Linguistics Olympiad (most challenging).
@@ 3.1 The effect of explicit operators in problems
Since so many of the mathematical operators in numeral structure are implicit (eg, in English we say ‘twenty three’ to mean ‘twenty + three’), our first experiment investigates how this implicit structure affects how models solve the problems. To do this, we standardize and convert the 10 existing linguistic number system problems to mathematical problems, and vary how explicit the operators are, as shown in Table 1.
First, we standardize all problems to control for model tokenization and task-external knowledge effects: we identify all meaningful morphemes, standardize all phonological changes, and replace them with dummy words as described in detail in Appendix A. This standardized version of each problem is what we call the implicit setting, since the mathematical operations are largely implicit, as they are in language. Taking these implicit problems as our baselines, we then make the operators explicit in three ways: 1) as the familiar mathematical operator symbols that perform the operation (eg, ‘+’ for addition), 2) as symbols that are unfamiliar for performing that operation, and 3) as whole words sampled from the tokenizer. A full example prompt with a puzzle in four variations is provided in Appendix B.
We present our results in Figure 1. In all cases, the presence of explicit operations with familiar symbols yields significant improvements over the default implicit condition (o1-mini performs at ceiling). In the multi-character setting (more linguistic), models perform better on average in the implicit condition than in the case with an explicit operator as an unfamiliar random word (vid. Figure 7). It is likely harder to differentiate between function words (operators) and number words (numerals) in such a setting — this finding is consistent with work that has shown human solvers also find a problem to be more difficult when the operator word is explicit but unfamiliar Derzhanski and Veneva (2018). Overall, our results demonstrate that it is difficult for models to reason about the abstract that linguistic quantities might contain operators, if the operators are not explicitly provided using familiar symbols.
@@ 3.2 Providing contextual information
Our first experiment showed that in the absence of problem-specific instructions, when given a linguistic-mathematical problem directly, LLMs struggle to solve it unless the operations are both explicit and familiar. This leaves open the question of whether providing additional problem-specific information would affect the model performance. We thus modulate the context of the problem in three different ways. We query the same four problem variants as described in Table 1, additionally providing the following contextual information:
@@
Figure 2: Language and base information only helps in the implicit case. Effect of adding language or numeral base information, plotted as a difference from the baseline values in Figure 1 for o1-mini. In cases with explicit operators, conflating overtly mathematical and linguistic information appears to confuse the models.
We compare these to the baseline results from Section 3.1 for o1-mini, presenting our results in Figure 2. In cases other than the implicit operator condition, the model seems to recognize the problem as requiring a more mathematical kind of reasoning, so providing linguistic information seems to confuse the model and average performance is worse. However, in the implicit operator (A B) condition, model performance improves significantly, perhaps because the setting of the problem is less overtly mathematical. In Figure 3, we show that providing information about the implicit reasoning needed is not as significant a boost as activating knowledge about the specific language.
Figure 3: Extra information improves performance on implicit problems (A B). Information about implicitness is helpful, but not as much as more direct information like the problem language. Error bars = standard error of the mean. 5 iterations / problem.
@@ 3.3 Ablations: constructed minimal-pair problems
In order to ensure that it is the difference in operators (as opposed to other features of the numeral system) that explains the models’ inability to solve these problems, we performed an ablation study to test whether models could handle other aspects of numeral construction and combination. Our experiment is inspired by the notion of a linguistic minimal pair, a pair of linguistic items that differ in exactly one meaningful element. We construct minimal pairs of simple, synthetic number system problems, where every element is the same except for one specific parameter that differs between two paired problems. We tested five major parameters of numeral systems, as described in Table 2.
@@
Parameter GPT-$3.5$ GPT-$4$ o1111-mini
Numeral system vs. not AB = fifty one | AB = big bird
Typed vs. glyph AB = fifty one | AB = 51
Order L $\rightarrow$ R vs. L $\leftarrow$ R AB = 51 | BA = 51
Additive vs. subtractive AB = 27 (20 + 7) | AB = 27 (30 - 3)
@@
Table 2: Minimal pair results: GPT-$4$ and o1-mini solve all paradigms, GPT-$3.5$-turbo struggles with numeral base and combination. Further data on testing all bases 4-19 linked in Table 5.
@@ Acknowledgments
This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence at Harvard University.
Appendix A Randomization strategies for task-external knowledge and tokenization handling
@@ Appendix B Multi-character-variable results from Exp 1
We provide an example of our four variations of the puzzle in Table 3. To query all four variants, we used the same prompt “Here is a puzzle. Can you solve it? Please output only the answer (in place of the ??) and nothing else!".
@@
(masaad $\times$ pagig) + masaad + opbob = 31 (masaad $\times$ pagig) + masaad + buylen = 26 $\vdots$ (ajssci $\times$ pagig) + (ajssci $\times$ kould) = 50 (innops $\times$ pagig) + innops + opbob = ??
@@
masaad pagig nge masaad opbob = 31 masaad pagig nge masaad buylen = 26 $\vdots$ ajssci pagig nge ajssci kould = 50 innops pagig nge innops opbob = ??
@@
(masaad $\beta$ pagig) $\alpha$ masaad $\alpha$ opbob = 31 (masaad $\beta$ pagig) $\alpha$ masaad $\alpha$ buylen = 26 $\vdots$ (ajssci $\beta$ pagig) $\alpha$ (ajssci $\beta$ kould) = 50 (innops $\beta$ pagig) $\alpha$ innops $\alpha$ opbob = ??
@@
(masaad hibcat pagig) xebrut masaad xebrut opbob = 31 (masaad hibcat pagig) xebrut masaad xebrut buylen = 26 $\vdots$ (ajssci hibcat pagig) xebrut (ajssci hibcat kould) = 50 (innops hibcat pagig) xebrut innops xebrut opbob = ??
Table 3: Example of four problem variants in the multi-character setting, corresponding to Drehu (IOL 2010) dataset problem in Figure 5.
@@ Appendix C Base experiment
In order to understand whether sufficiently advanced language models would show performance that was invariant to changes in the base, we conducted a more fine-grained minimal pair experiment into the effect of numeral base on problem performance. Here, the solver would see the Hindu-Arabic numerals corresponding to the English base-10 representation of the numbers, because the problem was presented in English. But the unknown symbols corresponded to the numbers as expressed in a different base, as shown in Figure 6.
@@
We conducted two different versions of this experiment. First, we mapped the unknown symbols to the single-character whitespaced A, B, C, and D tokens, as in Figure 6. In the second version, each of the four unknown symbols (A, B, C, D) was instead represented by a corresponding random token drawn from the tokenizer vocabulary, to ensure that the context of the specific tokens A, B, C, and D was not influencing our results.
We tested four increasingly sophisticated GPT models (GPT-$3.5$-turbo, GPT-$4$, GPT-$4$o, and o1111-mini) on both versions of the experiment and provide results in Table 5. GPT-4o and o1-mini solved all problems in both conditions, displaying performance that was robust to the base of the problem.
@@ Appendix D Table of languages
Base GPT-$3.5$-turbo GPT-$4$ GPT-$4$o o1111-mini
@@
Table 5: Base experiment results: GPT-$4$o / o1111-mini solve every problem, regardless of randomization
Figure 7: Both o1-mini and DeepSeek struggle with the explicit-unfamiliar condition (o1 shows negative improvement, DeepSeek shows 0%) in the multi-character setting. Error bars = standard error of the mean. 5 iterations / problem tested for 10 problems.
Figure 8: Results per language, (a) single-character (b) multi-character: performance varies significantly by problem and operator type.

### v2

@@ Abstract
Across languages, numeral systems vary widely in how they construct and combine numbers. While humans consistently learn to navigate this diversity, large language models (LLMs) struggle with linguistic-mathematical puzzles involving cross-linguistic numeral systems, which humans can learn to solve successfully. We investigate why this task is difficult for LLMs through a series of experiments that untangle the linguistic and mathematical aspects of numbers in language. Our experiments establish that models cannot consistently solve such problems unless the mathematical operations in the problems are explicitly marked using known symbols (++, $\times$, etc, as in “twenty + three”). In further ablation studies, we probe how individual parameters of numeral construction and combination affect performance. While humans use their linguistic understanding of numbers to make inferences about the implicit compositional structure of numerals, LLMs seem to lack this notion of implicit numeral structure. We conclude that the ability to flexibly infer compositional rules from implicit patterns in human-scale data remains an open challenge for current reasoning models.
@@
Antara Raaghavi Bhattacharya1, 2, Isabel Papadimitriou3, Kathryn Davidson2, David Alvarez-Melis1,3 1School of Engineering and Applied Sciences, Harvard University 2Department of Linguistics, Harvard University 3Kempner Institute for the Study of Natural and Artificial Intelligence at Harvard University Correspondence: antara@alumni.harvard.edu
@@ 1 Introduction
Figure 1: Making operators explicit significantly improves performance. Results for explicit operator experiments, for the single-character variable case (For the results on multi-character variables, see Appendix˜B Figure˜7). Making operators explicit shows performance improvement over the implicit condition, but this is only substantially and reliably the case when the operator is made explicit with a familiar symbol like “+”. Error bars denote standard error of the mean. 10 problems, 5 iterations per problem.
2 Background
2.1 Linguistic and cognitive connections
People acquire systems of number representation as part of learning language, and are consequently able to construct arbitrary numerals using the rules that they learn. Although the system of rules may be language-specific, the general framework of numeral construction and combination is a fundamental cognitive ability Hurford (1987); Feigenson et al. (2004). Performing mathematics in a symbolic sense requires explicit instruction (e.g. a child would not inherently know what + connotes), but once this symbolic meaning has been learned, people can generalize it to apply to any numbers Sarnecka et al. (2015).
Numeral operations in language can be marked both explicitly (e.g. und in German einundzwanzig) and implicitly (as in English twenty-one), with larger numerals often using a combination of implicit and explicit operations (five hundred and one = 5 $\times$] 100 + 1). Even when operations are implicit, people can understand and infer the cross-linguistic compositional structure of numerals Ionin and Matushansky (2006).
In a linguistics contest, a high-school student would not need to know any mathematical concepts beyond basic arithmetic to reason through number system problems and infer the rules needed to solve them. The challenge lies instead in whether models can learn and infer such rules from limited data — a characteristic capacity of humans acquiring language.
2.2 Mathematical ability in language models
Recent language models seem to display strong numerical understanding and processing abilities if presented with purely mathematical problems in standard formats Yang et al. (2024), particularly for small numbers and simple mathematical operations (of the kinds used in linguistics contest problems). Current reasoning models appear to perform well at arithmetic and algebra, math word problems Ahn et al. (2024), and difficult mathematical contest questions equivalent to advanced college-level math problems Fang et al. (2024); Chervonyi et al. (2025), although their problem-solving ability is sometimes inconsistent McCoy et al. (2023); Shojaee et al. (2025). If such models are unable to solve linguistic-mathematical problems involving much simpler mathematics, and introducing linguistic structure into the problem causes their reasoning ability to break down, this indicates limitations in the scope of their reasoning — models may be unable to apply their reasoning flexibly across domains in the ways that humans do.
@@ 3 Methods
Models. We used OpenAI o1-mini Jaech et al. (2024) and DeepSeek-R1-distill-Qwen-7B Guo et al. (2025) reasoning models to conduct our experiments, querying o1-mini via the API and running DeepSeek locally. All code and data used for our experiments are available at https://github.com/antara-raaghavi/multilingual-number-puzzles.
We additionally queried an instruction-tuned model (qwen-2-7b-vl-instruct) and a base model (llama-3.1-8B), both of which had an accuracy of 0 across all conditions that we test. These models almost always generated longer text answers without numbers rather than the simple numerical answer required, and were hence excluded from our analyses.
Data. We obtained data for linguistics olympiad problems from two publicly available datasets: LingOly Bean et al. (2024) and Linguini Sánchez et al. (2024), filtering both datasets for problems tagged as “number systems”. After filtering, we had 15 problems from the LingOly and 8 problems from the Linguini dataset. Not every problem in the dataset could be standardized in the ways that our experiments required. The entire dataset was thus manually evaluated for suitable problems, and 10 problems were chosen for evaluation, all in distinct languages (see Appendix˜D). These problems spanned a range of difficulty from the first round of the UK Linguistics Olympiad to the International Linguistics Olympiad (most challenging).
@@ 4.1 The effect of explicit operators in problems
Since so many of the mathematical operators in numeral structure are implicit (e.g. in English we say ‘twenty three’ to mean ‘twenty + three’), our first experiment investigates how this implicit structure affects how models solve the problems. To do this, we standardize and convert the 10 existing linguistic number system problems to mathematical problems, and vary how explicit the operators are, as shown in Table˜1.
First, we standardize all problems to control for model tokenization and task-external knowledge effects: we identify all meaningful morphemes, standardize all phonological changes, and replace them with dummy words as described in detail in Appendix˜A. This standardized version of each problem is what we call the implicit setting, since the mathematical operations are largely implicit, as they are in language. Taking these implicit problems as our baselines, we then make the operators explicit in three ways: 1) as the familiar mathematical operator symbols that perform the operation (e.g. ‘+’ for addition), 2) as symbols that are unfamiliar for performing that operation, and 3) as whole words sampled from the tokenizer. A full example prompt with a puzzle in four variations is provided in Appendix˜B.
We present our results in Figure˜1. In all cases, the presence of explicit operations with familiar symbols yields significant improvements over the default implicit condition (o1-mini performs at ceiling). In the multi-character setting (more linguistic), models perform better on average in the implicit condition than in the case with an explicit operator as an unfamiliar random word (vid. Figure˜7). It is likely harder to differentiate between function words (operators) and number words (numerals) in such a setting — this finding is consistent with work that has shown human solvers also find a problem to be more difficult when the operator word is explicit but unfamiliar Derzhanski and Veneva (2018). Overall, our results demonstrate that it is difficult for models to reason about the abstract idea that linguistic quantities might contain operators, if the operators are not explicitly provided using familiar symbols.
4.1.1 Error analysis
We observe some common patterns of error in the model responses. For the three problems which involved squares and cubes of numbers, when the operators were not explicit and familiar, o11-mini almost always responded by pattern-matching (e.g. providing another square/cube number) instead of solving the problem, as seen in Table˜2. o11-mini also reproduced a number given in the input question as the answer in several cases (11 for the multi-token condition, and 3 for the single-token condition, across 150 trials) when the operators were not explicit and familiar.
Condition Single Multi
explicit symbol 8 4
explicit random word 0 8
implicit 14 5
Table 2: Incorrect pattern-matched square / cube answers (out of 15 possible trials)
Further, when o11-mini answered a problem incorrectly, its responses were often inconsistent across the five trials of that problem. Notably, in 50% of single-character cases lacking explicit and familiar operators, all five responses were distinct and incorrect. This further shows that performance appears to depend on the presence of explicit operator cues; in their absence, o1-mini does not reliably solve the problem.
@@ 4.2 Providing contextual information
Our first experiment showed that in the absence of problem-specific instructions, when given a linguistic-mathematical problem directly, LLMs struggle to solve it unless the operations are both explicit and familiar. This leaves open the question of whether providing additional problem-specific information would affect the model performance. We thus modulate the context of the problem in three different ways. We query the same four problem variants as described in Table˜1, additionally providing the following contextual information:
@@
Figure 2: Language and base information only helps in the implicit case. Effect of adding language or numeral base information, plotted as a difference from the baseline values in Figure˜1 for o1-mini. In cases with explicit operators, conflating overtly mathematical and linguistic information appears to confuse the models.
We compare these to the baseline results from Section˜4.1 for o1-mini, presenting our results in Figure˜2. In cases other than the implicit operator condition, the model seems to recognize the problem as requiring a more mathematical kind of reasoning, so providing linguistic information seems to confuse the model and average performance is worse. However, in the implicit operator (A B) condition, model performance improves significantly, perhaps because the setting of the problem is less overtly mathematical. In Figure˜3, we show that providing information about the implicit reasoning needed is not as significant a boost as activating knowledge about the specific language.
Figure 3: Extra information improves performance on implicit problems (A B). Information about implicitness is helpful, but not as much as more direct information like the problem language. Error bars denote standard error of the mean. 5 iterations / problem.
@@ 4.3 Ablations: constructed minimal-pair problems
In order to ensure that it is the difference in operators (as opposed to other features of the numeral system) that explains the models’ inability to solve these problems, we performed an ablation study to test whether models could handle other aspects of numeral construction and combination. Our experiment is inspired by the notion of a linguistic minimal pair, a pair of linguistic items that differ in exactly one meaningful element. We construct minimal pairs of simple, synthetic number system problems, where every element is the same except for one specific parameter that differs between two paired problems. We tested five major parameters of numeral systems, as described in Table˜3.
@@
Parameter GPT-$3.5$ GPT-44 o11-mini
Numeral system vs. not
AB = fifty one | AB = big bird
Typed vs. glyph
AB = fifty one | AB = 51
Order L $\rightarrow$ R vs. L $\leftarrow$ R
AB = 51 | BA = 51
Additive vs. subtractive
AB = 27
(20 + 7)
AB = 27
(30 - 3)
@@
Table 3: Minimal pair results: GPT-44 and o1-mini solve all problem pairs, GPT-$3.5$-turbo struggles with numeral base and combination. DeepSeek-R1-distill-Qwen-7B does not produce an answer in the right format for most settings, so it is excluded from this table. Further data on testing all bases 4-19 linked in Table˜6.
@@ Acknowledgments
The authors gratefully thank Tom McCoy and Kaden Holladay for helpful discussions in the initial stages of this project. This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence at Harvard University.
Appendix A Randomization procedure for task-external knowledge and tokenization handling
@@ Appendix B Multi-character-variable results from Exp 1
We provide an example of our four variations of the puzzle in Table˜4. To query all four variants, we used the same prompt “Here is a puzzle. Can you solve it? Please output only the answer (in place of the ??) and nothing else!".
@@
(masaad $\times$ pagig) + masaad + opbob = 31
(masaad $\times$ pagig) + masaad + buylen = 26
$\vdots$
(ajssci $\times$ pagig) + (ajssci $\times$ kould) = 50
(innops $\times$ pagig) + innops + opbob = ??
@@
masaad pagig nge masaad opbob = 31
masaad pagig nge masaad buylen = 26
$\vdots$
ajssci pagig nge ajssci kould = 50
innops pagig nge innops opbob = ??
@@
(masaad $\beta$ pagig) $\alpha$ masaad $\alpha$ opbob = 31
(masaad $\beta$ pagig) $\alpha$ masaad $\alpha$ buylen = 26
$\vdots$
(ajssci $\beta$ pagig) $\alpha$ (ajssci $\beta$ kould) = 50
(innops $\beta$ pagig) $\alpha$ innops $\alpha$ opbob = ??
@@
(masaad hibcat pagig) xebrut masaad xebrut opbob = 31
(masaad hibcat pagig) xebrut masaad xebrut buylen = 26
$\vdots$
(ajssci hibcat pagig) xebrut (ajssci hibcat kould) = 50
(innops hibcat pagig) xebrut innops xebrut opbob = ??
Table 4: Example of four problem variants in the multi-character setting, corresponding to Drehu (IOL 2010) dataset problem in Figure˜5.
@@ Appendix C Base experiment
In order to understand whether sufficiently advanced language models would show performance that was invariant to changes in the base, we conducted a more fine-grained minimal pair experiment into the effect of numeral base on problem performance. Here, the solver would see the Hindu-Arabic numerals corresponding to the English base-10 representation of the numbers, because the problem was presented in English. But the unknown symbols corresponded to the numbers as expressed in a different base, as shown in Figure˜6.
@@
We conducted two different versions of this experiment. First, we mapped the unknown symbols to the single-character whitespaced A, B, C, and D tokens, as in Figure˜6. In the second version, each of the four unknown symbols (A, B, C, D) was instead represented by a corresponding random token drawn from the tokenizer vocabulary, to ensure that the context of the specific tokens A, B, C, and D was not influencing our results.
We tested four increasingly sophisticated GPT models (GPT-$3.5$-turbo, GPT-44, GPT-44o, and o11-mini) on both versions of the experiment and provide results in Table˜6. GPT-4o and o1-mini solved all problems in both conditions, displaying performance that was robust to the base of the problem.
@@ Appendix D Table of languages
Base GPT-$3.5$-turbo GPT-44 GPT-44o o11-mini
@@
Table 6: Base experiment results: GPT-44o / o11-mini solve every problem, regardless of randomization
Figure 7: Both o1-mini and DeepSeek struggle with the explicit-unfamiliar condition (o1 shows negative improvement, DeepSeek shows 0%) in the multi-character setting. Error bars denote standard error of the mean. 5 iterations / problem tested for 10 problems.
Figure 8: Results per language, (a) single-character (b) multi-character: performance varies significantly by problem and operator type. Note that Drehu (dhv) and Georgian (kat) are two of the easiest problems in our dataset: much of the difficulty for human solvers is in the phonological changes and unintuitive order that the numbers are presented in, both of which we standardize away for our controlled datasets. Without those parameters, the systems are straightforward vigesimal-decimal systems like French, which the models have almost certainly had exposure to.